OpenStudy (anonymous):

PLEASE HELP!!!! The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit

OpenStudy (anonymous):

Solve the equation for F and show all steps

OpenStudy (anonymous):

is that true? 100=212-32=180? no it is not true. There is a conversion factor of 5/9 you may obtain it using the known boiling point of water in celcius and fahrenheit 100=m(212-32) m=100/180=5/9 so C=5/9(F-32)

OpenStudy (anonymous):

F=9/5C+32

OpenStudy (anonymous):

\[F=\frac{9}{5}C+32\]

OpenStudy (anonymous):

ohhh, sorry about that, I didnt see that. I hade the 5/9 in there, but I guess it didnt enter.